OpenStudy (anonymous):

limx−>∞ of (√x2+4x−√x2−5x)

OpenStudy (anonymous):

**negative infinity

OpenStudy (turingtest):

\[\lim_{x \rightarrow -\infty}\sqrt{x^2+4x}-\sqrt{x^2-5x}\]?

OpenStudy (turingtest):

if so multiply it by\[\frac{\sqrt{x^2+4x}+\sqrt{x^2-5x}}{\sqrt{x^2+4x}+\sqrt{x^2-5x}}\]and simplify...

OpenStudy (anonymous):

i did that but im stuck at \[9x \div \sqrt{x^2+4x} +\sqrt{x^2-5x)}\]